OpenStudy (anonymous):

Identify the slope of the equation: y = 5x + 2 A. 4 B. 2 C. 7 D. 5

OpenStudy (binary3i):

y=mx+c where m is the slope find it

OpenStudy (anonymous):

i don't know .-.

OpenStudy (binary3i):

common try

OpenStudy (anonymous):

y=(5)x+2 y=mx+b do you see it now?

OpenStudy (anonymous):

is it c? .-.

OpenStudy (binary3i):

it is D

OpenStudy (anonymous):

because 5 is in the m spot